# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

### v1

@@ 5. Discussion
In Section 5.1 two of three usability questions and in Section 5.2 two of the three functionality questions are answered asked in the Methods. In Section 5.3 the last question of both studies will be discussed, which combines the perception of the online study and hands-on study. The resulting adaption features to improve the initial design are given in Section 5.4.
@@ 5.1.1. Are potential users able to use the system?
Yes, the survey showed in Q3, Q6, and Q8 (Figure 4) that most participants are able to use the system based on the video presentation (Answers: 17 yes, 3 no, 14 N/A). The high amount of N/A results were a result of including partially filled out surveys.
@@
In the open-ended questions, statements were made that the participants would like to have more information about the system to better evaluate the maintenance needed and the function of the system. These requests did not imply that the participant would change his or her mind after receiving the additional information. One participant stated “I wonder if the system can be used by people with intellectual disabilities due to their reduced cognition”. This underlines the importance of considering a variety of disabilities in the design process of such systems in order to make them accessible to a wide range of users. It highlights again the necessity to involve the community in the design process.
5.1.2. Which tasks are preferred so that the robot has the greatest benefit in everyday life?
@@
The categories of eating, drinking, environmental interactions, pick and place, personal hygiene, and miscellaneous tasks were found in the literature and are listed in Section 2.2. The participants’ responses from Section 4.1 were sorted into these categories that lead the distribution in Figure 8. The most requested task in the online survey that the robot should perform is assisting to eat or drink, followed by environmental interaction, which includes turning lights on and off, pressing buttons, opening doors, and pick and place objects. Picking up and moving objects was mentioned in different contexts such as moving heavy objects out of the way. Personal hygiene consisted of a variety of different tasks including doing makeup and hair, brushing teeth, and cleaning. Miscellaneous tasks included making phone calls and scratching themselves. These results can be corroborated with other findings (Chang and Park, 2003; Stanger et al., 1994; WHO, 2010). The variety of tasks expressed shows that participants were not biased by the tasks shown in the video.
[…]
This pie chart shows the classification of all mentioned tasks in relation to the presented system. 28 tasks were grouped into six categories containing the tasks mentioned by the participants and are shown in percentage. The largest category is nutrition with 52%, the second largest category is everyday tasks with 19%, the third is mobility with 11%, the fourth is two categories with 7% each described by the terms personal hygiene and well-being. Finally, tasks classified as communication are mentioned with 4%.
@@
Brushing one’s hair, applying makeup, or scratching are everyday tasks that require close human-robot interaction while ensuring the safety of the user. The shared control approach presented would require face and body recognition if the user is unable to move their head to a certain position. Ethical questions arise when personal data is recorded, including physical and physiological appearance. Current discussions in this area suggest that the user must be informed about the system functions and the personal data collected. This would allow the user to decide whether the privacy limitations or the benefits of the system outweigh (Fosch-Villaronga et al., 2020). The aim must be to present this information in simple language so that people with physical and cognitive disabilities can understand it and make decisions.
@@ 5.2.1. Would the prototype eye-tracking controller work in the controlled environment of the lab?
Initial testing showed that 70% of object selections would result in selection success when the dwell time trigger is set to 500 ms. Although, dwell times can be shortened to enhance selection success as shown in other literature (Holmqvist and Andersson, 2017), this could lead to an unintentional object selection as described in Section 2.3. An appropriate threshold must be explored with the prototype system.
@@
Accuracy issues are one of the main reasons why gaze was tracked outside of an AoI. Therefore, influences on AoI misses and dwell time were investigated in Section 4.2. Four dependencies were found:
@@
  • Object size: As shown in Figure 6, object size correlates significantly with dwell time. Small objects like keys are part of everyday activities. Excluding such objects from interaction, would lead to major drawbacks in participation in daily life. The available workspace is limited by the length of the robotic arm. This also limits the relative size of the object in the camera scene. With this limit, the resulting accuracy error can be calculated and checked with the desired interactable object size to estimate if this would cause problems. In this case, the accuracy of the eye-tracker is $\Delta_{deg}=0.6$°. With an effective workspace of $d=89.18$ cm for the used robot arm (inc., 2022), the maximum accuracy error is 9.34 mm ($\Delta=d*tan(\Delta_{deg})$). Since all selected objects were larger than 9.34 mm, it is assumed that the prototype would work.
@@
  • Geometry: The AoI of the fork could lead in a failed selection due to its width. Currently, there are two approaches in planning for estimating object outlines. Object recognition can be used to create bounding boxes, solving the issue due to the new geometry. Overlapping of bounding boxes may occur more often, leading to unintentional object selection and context misinterpretation. Object segmentation can lead to boundaries which fit to the object’s outlines. In this case the geometry of the object would be traced with high accuracy, solving the bounding box issue, leading back to the initial width issue.
@@
  • Gaze location towards the object: In the eating task example, participants tend to move their head to the left and right instead of up and down (see Figure 7). The resulting higher accuracy error was due to the glasses periphery. Considering physical disabilities of potential users, an immobility of the head cannot be excluded.
@@
  • User intention: Participants were instructed to focus their gaze on the selected object, but not on which region in detail. In the case of the bottle, the task was to fill the glass. The bottle needs to be opened first, which explains why the bottle cap was fixated, indicating the hand-eye coordination and intention behind the task as discussed in Section 2.1.
5.2.2. What are the technical difficulties with the design?
Ambiguous object selection can occur due to object overlap and was found in the case of the fork (see Figure 5). In addition to the accuracy challenges, robust object selection may be the most difficult challenge in implementing the prototype.
Task validation methods and emergency stops must be ensured, as found in Section 2.1. Approaches such as looking into the corner of the field of view could be challenging to interpret as an emergency stop because it is processed as the same data type as prolonged blinking in the eye-tracking software. Another approach describes the inclusion of gaze gestures (unique sequence of eye movements (Alsharif, 2017)). It needs to be verified whether such a cue would be able to perform for a person in an emergency situation. For task validation, the system needs to be enhanced with feedback methods to communicate the expected task by the robot.
@@ 5.3.1. Were the participants satisfied with the presented system?
In both studies, participants were mostly satisfied with the system. The challenges that were uncovered are summarized in Section 5.4.
5.3.2. Does participant perception vary greatly between the hands-on study and the online survey?
The unstructured in-person interview with participants in the hands-on study and the open-ended questions in the online survey, provided more diverse feedback on this shared control design. The results are consistent with and extend current literature in this area. It adds information for the design of gaze-controlled assistive robotic arms. Perceptions of the system do not vary but, highlight different aspects of the design.
Involving the stakeholder group yielded insights about everyday tasks and their daily life with their disability. The prior-experience group revealed potential drawbacks in the technical design process, such as concerns about safety, kinematic selection, and trajectory planning. Although the non-prior-experience group did not have in-depth feedback on certain aspects of the life with a disability or technology, they helped build a larger pool of participants to measure gaze behavior. In addition, their perception of the robot was more curious and at the same time hesitant than that of the prior-experience group, due to the novelty of robot interaction. Since robotic arms are rarely available to people with disabilities, it is reasonable to assume that in a real-world test, people with disabilities would respond in a initial robot interaction similarly to the non-prior-experience group.
5.4. Adaptation of the initial Shared Control design
Based on these results, the design can be refined to meet the participants’ expectations of the system. The next step is to automate the experimenter’s control task so that the robot can autonomously perform tasks selected by the user. The framework must have the following features to realize the system:
@@ 5.4. Adaptation of the initial Shared Control design
  • Gaze detection on objects: Due to the outcomes in Section 5.2, several approaches are followed to detect objects in the scene camera of the eye-tracking glasses. Unintentional selection and context misinterpretation will be estimated with the prototype.
@@
  • Task availability: The survey showed that the desired tasks were in line with the participants’ expectations. Eating was the most frequently mentioned task. However, handling a fork is a safety risk, when the robot manipulates the fork autonomously and has to be handled appropriately.
@@
  • Task selection: The expected prototype will implement objects associated with multiple tasks. Testing has shown that if the dwell time is set correctly and the algorithm is designed to wait for multiple objects to be selected, the task should be robustly selectable. Dwell times of 500 ms and less were found to be appropriate. As mentioned in Section 2.1, feedback strategies need to be evaluated and integrated into the control to ensure the user.
@@
  • Trajectory, grasping and collision avoidance: The robot has to perform the task by itself. In the current version, the robot’s trajectory was programmed and the objects were placed at the same location for each test. In a real-world scenario this would not meet the user’s expectations. Solutions were found as presented in Section 2.3) and depend on the approach chosen for object gaze detection.
@@
  • Adaptation: On behalf of the ambiguous results of the robot’s velocity perception, an adaptation setting will be included in the prototype so that users can adapt the system to their liking such as robot velocity.

### v2

@@ 5. Discussion
In Section 5.1, two of the three usability questions and in Section 5.2, two of the three functionality questions are answered. In Section 5.3, the last question of both studies will be discussed, which combines the perception of the online study and hands-on study. Feature improvements to the initial design are presented.
@@ 5.1.1. Are potential users able to use the system?
Yes, the survey showed in Q13(PSMD) and Q14(Fam./Fr. and Med. prof.) (Figure 4) that most of the participants are able to use the system based on the video presentation (Answers: 17 yes, 3 no, 14 N/A). The high amount of N/A results were the result of including partially filled out surveys.
@@
The limiting factors that prevent the system from being used every day were not stated. In the open-ended questions, participants stated that they would like more information. They also said that they would like more information about the system to better evaluate its maintenance and function. These requests did not imply that the participants would change their minds after receiving the additional information. One participant stated, “I wonder if the system can be used by people with intellectual disabilities due to their reduced cognition”. This highlights the importance of considering various impacts of disabilities during the design process to make such systems accessible to a wide range of users. It again highlights the necessity of involving the community in the design process.
Deeper issues, such as how the control works and whether wrong selections can occur due to natural gaze behavior, were not questioned by the participants, contrasting the results from the hands-on study. Section 5.3 shows participants’ concerns about interacting directly with the system, highlighting the benefit of dual investigations in robot perception.
5.1.2. Which tasks should be prioritized to maximize the robot’s usefulness in everyday life?
@@
The results, shown in Figure 5, the most requested tasks are assisting while eating or drinking, followed by environmental interaction, which includes turning lights on and off, pressing buttons, opening doors, and picking and placing objects. Personal hygiene consisted of a variety of different tasks, including doing makeup and hair, brushing teeth, and cleaning. Miscellaneous tasks included making phone calls and scratching themselves.
@@
Brushing one’s hair, applying makeup, or scratching are daily tasks that require close human-robot interaction while ensuring the safety of the user. The shared control approach presented would require face and body recognition if the user is unable to move their head to a certain position. Ethical questions arise when personal data is recorded, including physical and physiological appearance. Current discussions in this area suggest that the user must be informed about the system functions and the personal data collected. This would allow the user to decide whether the privacy limitations or the benefits of the system outweigh (Fosch-Villaronga et al., 2020). The aim must be to present this information in simple language so that people with physical and cognitive disabilities can understand it and make decisions.
The prioritized tasks and the tasks realized from the literature included in Section 2.3 were summarized. A comparison of the results with the findings in the literature shows partial agreement. Table 4 compares findings from the literature with our results. In the ”Prioritization” column, we summarized the works of Beer et al., Nanavati et al., Petrich et al., Smarr et al., Stanger et al., Chang et al., and Jardón et al. (Beer and Akanksha Prakash, 2011; Petrich et al., 2022; Nanavati et al., 2024b; Smarr et al., 2012; Stanger et al., 1994; Chang and Park, 2003; Jardón et al., 2012). These works present or review task prioritization in activities of daily living (ADLs), involving either elderly people or people with disabilities. It is important to acknowledge that these works include varying participant groups, as well as varying robot types. To the best of the authors’ knowledge, no review or work solely focuses on PSMD and robotic arms. The ”Integration” column summarizes reviews and individual studies concerning the execution of tasks with assistive robots. It summarizes the outcomes of the following works, which were also presented in Section 2.3 (Nanavati et al., 2024b; Hagengruber et al., 2025; Dragan and Srinivasa, 2013; Huang and Mutlu, 2016; Zhang et al., 2025; Bien et al., 2003; Canal et al., 2021; Bhattacharjee et al., 2020). Some tasks were excluded from the literature when they could not be performed with a robotic arm (e.g., getting around or into bed, going out, toileting, managing money, leisure and recreation, and shopping), as well as outdated tasks (e.g., removing paper from printers and fax machines).
Table 4. Overview of tasks mentioned by participants and representation within literature. Tasks mentioned by users are presented to their according categories in Figure 5. Fit describes if the task can be accomplished without additional assistive devices to a robotic arm (triangle shows partial accomplishment). The comparison is done between our results, task prioritization found in other works and integrated tasks in existing works.
The table presents an overview of daily-life task categories, specific tasks within each category, and how these tasks are represented in the collected user responses and in prior literature on assistive robotic arms. Categories include Eating, Drinking, Environmental Interaction, Personal Hygiene, and Miscellaneous Tasks. For each task, the table indicates whether it can be performed by a robotic arm without additional assistive devices, and compares three aspects: how frequently the task was mentioned by participants in the hands-on study, how strongly participants in related literature studies preferred the task, and how strong the research interest in realizing the task in existing assistive robotic systems is. Eating and drinking are both considered feasible with a robotic arm alone and are frequently mentioned by users, commonly prioritized in literature, and often integrated in existing systems. Environmental interaction tasks include preparing meals, using switches and buttons, opening doors and drawers, pick-and-place actions, and carrying objects. All of these tasks can be accomplished with a robotic arm alone. All tasks except preparing meals are mentioned by participants in this study. They are highly prioritized by participants in other studies, and commonly integrated in existing systems, with Switches, Buttons and carrying objects showing slightly less frequent integration. Personal hygiene tasks include makeup or shaving, combing hair, brushing teeth, dressing, and bathing or grooming. All are feasible with a robotic arm alone, except bathing and grooming. Bathing and grooming are only partially feasible without additional assistive devices. They mentioned by participants in this study except Dressing and Bathing which also represents the results from other literature. They are less commonly integrated into existing systems. Miscellaneous tasks include using phones or electronics, taking medicine, doing household chores, and scratching. Household chores are feasible, mentioned in this study and literature and frequently integrated in existing systems. Using electronics and scratching are only partially feasible without additional devices. Taking medicine is feasible but less frequently mentioned and less often integrated in existing systems.
Category Tasks Fit Our results Prioritization Integration
Eating Eating $\checkmark$ $\nearrow$ $\rightarrow$ $\uparrow$
Drinking Drinking $\checkmark$ $\nearrow$ $\nearrow$ $\uparrow$
Environmental Preparing meals $\checkmark$ $\uparrow$ $\uparrow$
Interaction Switches, Buttons $\checkmark$ $\nearrow$ $\uparrow$ $\rightarrow$
Doors, drawers $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
Pick and place $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
Carrying objects $\checkmark$ $\nearrow$ $\uparrow$ $\nearrow$
Personal Makeup, shaving $\checkmark$ $\nearrow$ $\uparrow$ $\rightarrow$
hygiene Combing hair $\checkmark$ $\nearrow$ $\nearrow$ $\searrow$
Brushing teeth $\checkmark$ $\nearrow$ $\nearrow$ $\searrow$
Dressing $\checkmark$ $\rightarrow$ $\rightarrow$
Bathing, Grooming $\triangle$ $\rightarrow$ $\nearrow$
Miscellaneous Using phone, electronics $\triangle$ $\nearrow$ $\rightarrow$ $\rightarrow$
tasks Taking medicine $\checkmark$ $\rightarrow$ $\searrow$
House choirs $\checkmark$ $\nearrow$ $\uparrow$ $\uparrow$
Scratching $\triangle$ $\nearrow$ $\nearrow$ $\searrow$
Tasks such as preparing meals, dressing, bathing, grooming, and taking medicine were not mentioned by the participants. This may be because these tasks are comparatively less important to end users, as found in the literature. In contrast, preparing meals is highly prioritized and researched in other studies. Depending on their hobbies or the type of personal assistance they receive, this task may not have been favored by the small group of participants. The insight into the eating task is interesting. While most participants stated that they strongly preferred to accomplish this task independently, Nanavati et al. found that, compared to other ADLs and IADLs, this task was the least prioritized (Nanavati et al., 2024b). However, other sources, such as Petrich et al. and Stanger et al., found it to be highly prioritized (Stanger et al., 1994; Petrich et al., 2022). This discrepancy may be due to the fact that Nanavati et al.’s work reviewed different populations, including people with visual impairments and the elderly, who may be able to eat without assistance.
Comparing works that address these tasks reveals gaps in interacting with switches and buttons, personal hygiene, scratching, and using electronics. Concerns about user safety and privacy regulations might influence the low number of works related to personal hygiene and scratching. In contrast, using electronics or phones is becoming more convenient with smart home applications and hands-free technology. In conclusion, the chosen task for the presentation aligns well with the literature. More focus should be placed on assisting people with personal hygiene tasks. The variety of tasks performed demonstrates that the participants were not influenced by the tasks shown in the video.
@@ 5.2.1. Would the prototype eye-tracking controller work in the controlled environment of the lab?
Initial testing showed that 70% of object selections would result in selection success when the dwell time trigger is set to 500 ms. Although dwell times can be shortened to enhance selection success, as shown in other literature (Holmqvist and Andersson, 2017), this could lead to an unintentional object selection, as described in Section 2.1. An appropriate threshold must be explored with the prototype system.
@@
Accuracy issues are one of the main reasons why gaze was tracked outside of an AoI. Therefore, the influences on AoI misses and dwell time were investigated in Section 4.2. Four dependencies were found:
@@
  • Object size: As shown in Figure 7, object size correlates significantly with dwell time. Small objects, like keys, are part of everyday activities. Excluding such objects from interaction would lead to major drawbacks in participation in daily life. The available workspace is limited by the length of the robotic arm. This also limits the relative size of the object in the camera scene. With this limit, the resulting accuracy error can be calculated and checked with the desired interactive object size to estimate if this would cause problems. In this case, the accuracy of the eye-tracker is $\Delta_{deg}=0.6$°. With an effective workspace of $d=89.18$ cm for the used robot arm (inc., 2022), the maximum accuracy error is 9.34 mm ($\Delta=d*tan(\Delta_{deg})$). Since all selected objects were larger than 9.34 mm, it is assumed that the prototype would work.
@@
  • Geometry: The AoI of the fork could lead to a failed selection due to its width. Currently, there are two approaches in planning for estimating object outline. Object recognition can be used to create bounding boxes, solving the issue due to the new geometry. Overlapping of bounding boxes may occur more often, leading to unintentional object selection and context misinterpretation. Object segmentation can lead to boundaries which fit the object’s outline. In this case, the geometry of the object would be traced with high accuracy, solving the bounding box issue, leading back to the initial width issue.
@@
  • Gaze location towards the object: In the eating task example, participants tend to move their heads to the left and right instead of up and down (see Figure 8). The resulting higher accuracy error was due to the glasses periphery. Considering physical disabilities of potential users, an immobility of the head cannot be excluded.
@@
  • User intention: Participants were instructed to focus their gaze on the selected object, but not on which region in detail. In the case of the bottle, the task was to fill the glass. The bottle needs to be opened first, which explains why the bottle cap was fixated, indicating the hand-eye coordination and intention behind the task, as discussed in Section 2.1.
5.2.2. What technical difficulties arise from the design?
Ambiguous object selection can occur due to object overlap and was found in the case of the fork (see Figure 6). In addition to the accuracy challenges, robust object selection may be the most difficult challenge in implementing the prototype. Other researchers have addressed this issue using 3D gaze tracking to minimize localization errors in approaches involving eye-tracking glasses (Cao et al., 2025; Yang et al., 2021). As discussed in Section 2.1, user intent may lead to more robust selection methods by involving multiple objects to estimate task intent. These objects may vary in geometry and size, which reduces the impact of incorrect fixations. Regarding current approaches using object detection, overlap in bounding boxes of closely arranged objects may lead to larger issues than those found in the area of interest (AoI) evaluation. The AoIs were fitted to the outlines of the objects, similar to object segmentation. However, bounding boxes have larger areas, which leads to ambiguity when they overlap.
Section 2.2 presents technical challenges, level of autonomy, and user preferences as factors influencing user acceptance. Within these categories, we found indications from the survey and hands-on study of technical difficulties.
After implementing the suggested control method, the tasks will be selectable as planned. However, as Hagengruber et al. (Hagengruber et al., 2025) point out, accomplishing subsequent tasks is necessary to realize benefits for everyday usage. With this in mind, objects could indicate multiple tasks. Solutions integrating machine learning methods to identify gaze patterns, as presented by Yang et al., Jain et al., and Belardinelli, can reduce misinterpretation (Yang et al., 2021; Jain and Argall, 2019; Belardinelli, 2024).
In the WoZ design, the robot was presented with a high level of autonomy. As discussed in Section 2.2, it is questionable whether participants will accept the system as intended. In the hands-on study, participants vocalized their safety concerns, especially in the group with no prior robot experience. Arbitration methods can improve perception. This is especially important, since robot perception changes with regular use, and the level of autonomy can be reduced until the user is confident enough to operate the robot with a higher level of autonomy. Additionally, settings should be integrated that allow users to adjust the robot’s velocity to their preferred level.
@@ 5.3.1. Were the participants satisfied with the presented system?
In both studies, the participants were mostly satisfied with the system. We found in both studies suggestions for improvement. The design can be refined to meet the participants’ expectations of the system. The framework must have the following features to realize the system:
@@
  • Gaze detection on objects: Due to the outcomes in Section 5.2, several approaches are followed to detect objects in the scene camera of the eye-tracking glasses. The prototype will estimate unintentional selection and context misinterpretation.
@@
  • Task availability: The survey showed that the desired tasks were in line with the participants’ expectations. Eating was the most frequently mentioned task. However, handling a fork is a safety risk when the robot manipulates the fork autonomously, and has to be handled appropriately.
@@
  • Task selection: The expected prototype will implement objects associated with multiple tasks. Testing has shown that if the dwell time is set correctly and the algorithm is designed to wait for multiple objects to be selected, the task should be robustly selectable. Dwell times of 500 ms and less were found to be appropriate. As mentioned in Section 2.1, feedback strategies need to be evaluated and integrated into the control to ensure the user. Strategies have to be evaluated if user intent methods assist in solving issues selecting between multiple tasks.
@@
  • Trajectory, grasping, and collision avoidance: The robot has to perform the task by itself. In the current version, the robot’s trajectory was programmed, and the objects were placed at the same locations for each test. In a real-world scenario, this would not meet the user’s expectations. Solutions were found, as presented in Section 2.1), and depend on the approach chosen for object gaze detection.
@@
  • Adaptation: Due to the ambiguous results of the robot’s velocity perception, settings will be included in the prototype so that users can adapt the system to their liking, such as robot velocity.
In future work, the experimenter’s control has to be automatized so that the robot can autonomously perform tasks selected by the user.
5.3.2. Does perception vary between the hands-on study and the online survey?
The unstructured, in-person interviews with participants in the hands-on study and the open-ended questions in the online survey provided more diverse feedback on this shared control design. The results are consistent with, and extend, current literature in this area. It adds information for the design of gaze-controlled assistive robotic arms. Perceptions of the system do not vary, but highlight different aspects of the design.
Involving the stakeholder group yielded insights about daily tasks and their daily life with their disability. The prior-experience group revealed potential drawbacks in the technical design process, such as concerns about safety, kinematic selection, and trajectory planning. Although the non-prior-experience group did not have in-depth feedback on certain aspects of life with a disability or technology, they helped build a larger pool of participants to measure gaze behavior. In addition, their perception of the robot was more curious and, at the same time, hesitant than that of the prior-experience group due to the novelty of robot interaction. Since robotic arms are rarely available to PSMD, it is reasonable to assume that, in a real-world test, PSMD would respond in an initial robot interaction similarly to the non-prior experience group.